What is minimum of f(x, y) = x² + y2 subject to 3x + 4y = 25
Answer #1

sobre Now floxg) = x + y² subject to 3x + 4y = 25 f (x, y) = scty2 , 3x + 4y= 25 - (Y) = 0 If (8,y) = xx (XY) (fxrby> = x&xid